Circular No. 8472

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005N, 2005O, AND 2005P Further to IAUC 8470, J. Burket and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on Jan. 21.55 (at mag 18.1) and 22.54 UT (mag 18.0). SN 2005P is located at α = 14^{h}06^{m}34^{s}.01, δ = −5^{o}27′42′′.6 (equinox 2000.0), which is 12′′.6 west and 31′′.7 south of the nucleus of NGC 5468. A KAIT image taken on 2004 July 8.21 showed nothing at this position (limiting mag 19.5).

B. Schmidt and M. Salvo, Australian National University, report that a spectrogram (range 420–1000 nm) of SN 2005O (cf. IAUC 8471), obtained by P. Wood on Jan. 21.70 UT with the ANU 2.3-m telescope (+ DBS), reveals it to be a type-Ib supernova near maximum light. The relatively red spectrum is dominated by He I lines and the Ca II infrared triplet.

S. Taubenberger and A. Pastorello, Max-Planck-Institut für Astrophysik, Garching, on behalf of the European RTN collaboration, report that inspection of a low-S/N spectrogram of SN 2005N (cf. IAUC 8470), taken on Jan. 22.21 UT by M. Alises with the Calar Alto 2.2-m telescope (+ CAFOS; range 330–880 nm), shows it to be a type-Ib/c supernova in the nebular phase, ∼ 4 months after the explosion. The spectrum is characterized by strong emission lines of [O I]; features due to [Ca II], Mg I], [Fe II], Na I, O I, and Ca II are also visible.

SUPERNOVA 2001em IN UGC 11794 C. J. Stockdale and B. Kaster, Marquette University; L. O. Sjouwer-

man and M. P. Rupen, National Radio Astronomy Observatory; I. Marti-Vidal and J. M. Marcaide, University of Valencia; S. D. Van Dyk, Spitzer Science Center, California Institute of Technology; K. W. Weiler, Naval Research Laboratory; B. Paczynski, Princeton University; and N. Panagia, European Space Agency and Space Telescope Science Institute, report a test of the proposal by Paczynski (2001, Acta Astron. 51, 1) and by Granot and Loeb (2003, Ap.J. 593, L81) that type-Ib/c supernovae may produce latetime radio jet emission. A Very Long Baseline Array measurement on 2004 July 1 of the oldest known radio-emitting type-Ib/c supernova, SN 2001em, at age ∼ 1020 days after explosion, shows no extended radio emission to a 3σ limit of 0.3 mJy/beam. The detected radio emission at 8.4 GHz of 1.8 ± 0.2 mJy from SN 2001em appears unresolved at a resolution of 0′′.0019 × 0′′.0008, corresponding to a radius of < 0.4 ± 0.2 pc at a distance of 90 Mpc or an average expansion speed for any detectable radio emitting material of < 150000 km/s (< 0.5c).

2005 January 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green